4. Suppose the value of your portfolio will either double or fall with equal probability in any particular year.

a. What is the expected value of the portfolio after one year?

b. What is the expected value of the arithmetic average return on the portfolio?

c. What is the expected value of the geometric average rate of return?

d. In light of the answer to part (a), would you consider the arithmetic or geometric return more informative for performance evaluation? 
Answer #1

1.
=0.5*2+0.5*0.5
=1.25

2.
=(100%-50%)/2=25%

3.
=((1+100%)*(1-50%))^(1/2)-1=0.000%

4,
Geometric return